A corporation needs you to calculate the carrying value (book value) of a piece of PP&E at the end of the second year.  

Depreciation Method Used => Double-Declining-Balance Method

  • Life of PP&E Asset => 8

  • Invoice Price of PP&E Asset => 121000

  • Setup Charges of PP&E Asset => 54000

  • Salvage Value of PP&E Asset => 9000
Answer #1

Answer : Carrying value of PPE at the end of two years = 98437.

Solution -

Depreciation rate = [(100/8)×2]

= 12.5×2

= 25

That means, rate of depreciation = 25%

Depreciation for 1st year = 175000×25% = 43750

Depreciation for 2nd year = (175000-43750)×25% = 32813

Therefore,

Carrying value of PPE at the end of two years = (121000+54000) - (43750+32813)

= 175000 - 76563

= 98437.